OpenStudy (anonymous):

y = –3(x + 7)^2 – 10. Y=x^2. How is this equation shifted on a graph? Up, down open or closed, narrow etc...

OpenStudy (anonymous):

y = –3(x + 7)^2 – 10. Y=-3(X^2+14X+49)-10 Y=-3X^2-42X-147-10 Y=-3X^2-42X- 157

OpenStudy (anonymous):

okay. is there more to it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay, can you explain how it's shifted plz.

OpenStudy (precal):